# The MEMENTO code for modelling of macroscopic melt motion in fusion devices

### v1

@@ 1 Introduction
Power handling and plasma-facing component (PFC) longevity constitute major technological obstacles on the way to efficient operation of fusion reactors [1]. The melting of metallic armor under high energy-density transient events is the main threat to PFC integrity, since liquid metal displaced by plasma-induced or external forces can cause large-scale surface deformations and even modify the castellated PFC structure. Since contemporary fusion devices cannot achieve the plasma parameters of future machines, such as ITER [1] & DEMO [2], predictive modelling of thermal response and melt motion is crucial in guiding reactor designs and choices of safe operational space.
@@
The physics model and treatment of plasma-related effects have been guided by empirical evidence from a series of dedicated tokamak experiments [8, 9, 10, 11, 12, 13]. The numerical implementation of the model in the eponymous code [4, 5, 6] has been correspondingly carried out in an additive fashion. Moreover, as validation against experiments proved successful, the use of the tool for predictive modelling for fusion reactors became essential. Such applications pose additional challenges due to the necessity to simulate intricate PFC layouts as well as complex plasma-wetted geometries. Furthermore, simulations of at least an entire plasma-facing unit (monoblock, castellation) are necessary, where it is essential to resolve the thin localized melt pools on timescales that include long power loading combined with short quasi-periodic bursts. The lack of a priori knowledge of the PFC geometries and wetting details as well as the spatiotemporal scales to be tackled inhibited the development of efficient numerical solutions. The original implementation of the MEMOS-U physics model was computationally inefficient and lacked flexibility in simulations of complex domains and complicated scenarios. This served as the primary incentive behind the development of a new implementation of the physics model in a modern computational tool - MEMENTO (MEtallic Melt Evolution in Next-step TOkamaks).
In this work, an in-depth overview of MEMENTO is presented, including a detailed description of the adopted numerical schemes. The MEMENTO code relies on AMReX, a software framework that provides functionality for massively parallel codes and adaptive meshing techniques [22, 23]. Numerical benchmarks of the solvers of MEMENTO have been presented elsewhere [24]. Here, we discuss the code structure and provide a simulation flowchart as well as examples of the MEMENTO capabilities in the modelling of fusion-relevant melting events.
@@ 2 Model equations
In what follows, for the sake of completeness, we shall recap the set of equations as well as introduce all the physical quantities and thermophysical properties involved:
@@
In the above, the liquid column height is defined as , with and denoting the evolving surface of the liquid-solid interface and the evolving metal-plasma interface (free surface) respectively, as sketched in Fig.1. The SW equations (1)-(2) describe the evolution of the film thickness and the depth-averaged velocity ; is the rate of change of due to vaporization, is a unit vector tangential to the free surface, the bulk current density whose depth-averaged value is defined as
the external magnetic field, the mass density, the surface tension, the drag force and the ambient pressure applied by the plasma and vapour at the surface. In the heat equation (3); is the specific isobaric heat capacity, the thermal conductivity, the volumetric heat sources, any bound of the computational domain for which Eq.(3) is solved, the vector unit normal and the surface heat flux. In the EM equations (4)-(5) which enforce the conservation of charge; is the electrical resistivity, an auxiliary potential and the emitted current density. Three surface boundaries are distinguished; is the free surface with the unit vector normal , is any electrically insulated surface with the unit vector normal and is any grounded surface.
@@ 3 Code structure
MEMENTO relies on the AMReX framework [22, 23] to generate and maintain an adaptive non-uniform grid. The non-uniform grid allows MEMENTO to perform highly resolved calculations near the free surface and to save computational time by introducing a coarser grid far from it. Each distinct grid resolution defines an AMReX level, with the free surface always corresponding to the finest level. The coarsest level is always designated as “level 0”. The terms higher and finer will be used interchangeably when referring to AMReX levels. The ratio between grid resolutions of two different levels is referred to as the refinement ratio; it is defined in the input and has to be a power of two in all MEMENTO simulations. Fig.2 shows an example of the cross-section of a simulated sample and the employed three-level grid. Owing to the adaptive nature of the mesh, the grid has changed between the 5 and 11 ms instants due to the build-up of a strong surface deformation. To complement the non-uniform spatial mesh, sub-cycling in time can be enabled so that higher levels are advanced with a smaller time-step. The ratio between the time-steps of different levels is equal to their refinement ratio.
@@ 4.1 Implementation
The parameters and are controlled by the input and serve as ad-hoc solutions to possible numerical issues by capping the viscous damping and the Marangoni drive to a minimum effective fluid height. The designation refers to the partially updated height obtained by solving for the position of the free surface in time step but not for the solid-liquid interface, i.e. . It is noted that the explicit formulation of the surface tension introduces a stringent stability criterion for the time-step that scales poorly with the tangential discretizations and [30]. Moreover, the Lorentz force, , is a variable that has to be passed from the electrostatics solver with the following interpolations employed
@@ 5 Heat transfer solver
The velocity field employed by the heat solver, , is a 3D vector, whereas the depth-averaged velocity from the SW solver is a 2D vector. Reconstruction of from is based on consistent with the SW assumption of . For ease of notation, in the following, we will simply express as .
@@ 5.2.1 Implementation
The time-step indicator ∗ has the same meaning as the one discussed in the explicit heat solver, see section 5.1.1.
@@ 7 MEMENTO simulations
Possible simulated geometries are categorized in two groups; slab and complex. In slab geometries the initial free surface lies at a constant height, , while in complex geometries the free surface is a function of position, . Examples of complex geometries include free surfaces that are initially sloped, curved or step-like. The free surface is wetted by default, whereas the wetting of other surface boundaries is controlled by the input. Cooling is possible from all surfaces, but mass-flux due to vaporization is only defined for the free surface. The sample can include a heat sink in the solid which emulates a cooling pipe, as the one present in ITER monoblocks. Finally, it is possible to include regions where the thermal and electrical conductivities are set to zero, to emulate gaps in the domain.
@@ 8 Re-gridding, re-evaluation and synchronization
Re-gridding refers to the modification of the grid used by the heat and electrostatics solvers to comply with the deformation of the free surface. The goal is to keep the finely resolved region to the minimum size required at each time-instant in order to reduce the computational cost. In the current version of MEMENTO, there are three defined criteria that can trigger re-gridding to a higher or lower level. Those are the distance of a point from the free surface, the surface deformation rate and the material phase at a certain point. The values of the distance from the free surface and the deformation rate that trigger re-gridding are defined in the input. Note that the surface deformation rate is a function of and not . Hence, if the value triggers re-gridding then all the points for which are re-gridded. The third criterion sets all the points in the liquid phase at the highest level, thus the re-gridding is triggered as the melt front propagates. By default these criteria are checked at every time-step while different frequencies can be chosen via input. The default option is to check for re-gridding at every time step. In Fig.2, at 5 ms re-gridding has been triggered only due to the distance from the free surface while at 11 ms also the surface deformation rate triggers re-gridding.
@@ 9 Solver benchmarking and comparison with experiments
Before illustrating the MEMENTO capabilities through the simulation of a tokamak experiment, we briefly recap the code benchmarking reported in detail elsewhere [24]. The fluid shallow water solver has been compared with the augmented solver of the GeoClaw software [32, 33, 34, 35]. The accuracy of the explicit heat solver has been checked against analytical solutions of a two-region Neumann problem [36]. The implicit heat solver has been benchmarked utilizing the explicit heat solver as reference code. Furthermore, the MEMENTO implementation of a cooling pipe, based on simplifications suggested in Ref.[37], has been compared to the results of ANSYS simulations [38] and to the results of simulations with a dedicated solver [39]. Finally, the electrostatics solver has been validated against simulations with the COMSOL Multiphysics® software. The numerical inaccuracies were always found to be negligible compared to typical uncertainties of the MEMENTO input, primarily the experimental heat fluxes which often feature uncertainties in excess of 20 [3, 5, 6, 13].
@@ 10 Example simulation of a tokamak experiment
Given the above, we have selected the simulation of a particular recent tokamak experiment in order to demonstrate the capabilities of MEMENTO. This experiment, featuring several momentum sources, a non-trivial initial free surface, a non-slab sample geometry and a particular plasma wetting pattern, exemplifies a complex scenario which MEMENTO has been designed to model.
@@ 10.1 Experimental scenario
Here we model the exposure of an iridium (Ir) sample to ASDEX Upgrade (AUG) tokamak discharges [13]. The main physics aspects have been reported in Ref.[13]. In what follows, we focus on numerical considerations concerning the simulation set-up and post-processing.
@@ 10.3 Simulation outcome
Finally, the deformation profile is a key output of the simulation, since it can be directly compared to the experiment. Such a comparison is presented in Fig.10 revealing that MEMENTO is capable to reproduce the post-mortem evidence both qualitatively and quantitatively. Note that the shift between the MEMENTO deformation profile and the observations is well within the experimental uncertainties on the position of the heat-flux on the tile ( cm). For more detailed discussions on the melt dynamics and the final deformation, we address the reader to Ref.[13]. Once again, it is pointed out that deformation is accumulated from all discrete melting events and that resolidification takes place in between melting.
@@ 11 Summary and discussion
We shall briefly recap the status of the MEMENTO solvers to point out the necessity for future updates. We shall also discuss inherent limitations of the code in specific fusion-relevant applications, due to the underlying shallow water assumption.
@@
The fluid solver provides the solution to the SW equations. Depending on the scenario, to address stability issues inherent to the shallow water model, it might require truncation of the melt depth column to some minimal values, set by the input. The introduction of such melt depth thresholds can be entirely avoided or set to negligible values by including the surface tension in the simulations; an optimal choice both from a physics and a numerical point of view. However, in scenarios where the melt pool is created at or moved to an edge, the SW approximation cannot capture the 3D nature of the resulting pool curvature and the details of the corner wetting, which is a limitation of the physics model rather than of the implementation. In case the scenario features an insignificant forcing of the pool off the edge, then an impermeable wall can provide an adequate solution. Such a solution was employed in the modelling of a recent AUG melting experiment, where all features of the experimental surface deformation were successfully recovered in the MEMENTO runs [14]. Future optimization of this module primarily concerns the surface tension implementation. In particular, other numerical schemes will be investigated in an attempt to improve the flexibility with respect to the choice of grid resolution.
@@
Finally, we stress again that the shallow water approximation constitutes the major reason for MEMENTO’s ability to reliably model fusion-relevant melting scenarios where meter-long plasma-wetted areas must be simulated while resolving micrometer-deep melt pools. Simultaneously, the shallow water assumption also imposes constraints on certain types of scenarios. The associated limitations are mainly confined to two cases. The first limitation concerns the presence of near degree slopes in the component geometry, where projection to the shallow water plane might lead to ambiguous results. This is not the case with modest slopes as demonstrated in the modelling example of section 10, where MEMENTO simulations quantitatively matched all experimental features of the melting events taking place on a sloped sample. The second limitation concerns cases with excessive melting and forcing that result in drastic material excavations and a spatio-temporally varying SW plane projection. Examples of such experimental situations have been reported in Refs.[13, 14].
Since the shallow water originating limitations mainly concern the aforementioned two cases, it can be stated that the majority of complex fusion-relevant melting scenarios can be attacked with the MEMENTO simulations. Moreover, in order to facilitate the self-consistent modelling of small-scale melt dynamics and melt flow stability in fusion reactors, the MEMENTO simulation outcome can be utilized in specialized ANSYS CFD set-ups. Such a work-flow enables the faithful reproduction of the macroscopic picture (large wetted areas, three-dimensional heat transfer, long time thermal response, large scale melt dynamics) responsible for the eventual melt pool depths and velocities , while allowing for full detail simulations of a small domain of interest where and are imposed through boundary and initial conditions [42]. After the necessary tests against experimental evidence from melting events in the JET tokamak [42], this work-flow has been employed for predictive modelling of melt splashing in ITER [43] and will be utilized in predictive studies for DEMO.
@@ Appendix A Simulation flowchart
The flowchart of a MEMENTO simulation is shown in Fig.1. In short, a simulation starts by reading the input and performing the necessary initializations. The free surface is generated at the specified position, the initial temperature is set, the melt thickness is initialized to zero and the material properties are calculated. If sub-cycling is not enabled, the fluid solver is called first, followed by the heat solver and the electrostatics solver. If sub-cycling is enabled, the order of operations changes. The heat equation is first solved on all levels except the highest. Then, the SW, heat and electrostatics solvers are called in sequence and repeatedly according to the time-step division by the refinement ratio. Once all solvers are updated, the code checks an input-defined frequency to determine whether output is to be generated in the current time step. Finally, the termination conditions are checked to see whether the simulation has reached its end. If not, time is incremented and the calculations are repeated.

### v2

@@ 1 Introduction
Power handling and plasma-facing component (PFC) longevity constitute major technological obstacles on the way to efficient operation of fusion reactors [1]. The melting of metallic armor under high energy-density transient events is the main threat to PFC integrity, since liquid metal displaced by plasma-induced or external forces can cause large-scale surface deformations and even modify the castellated PFC structure. Since contemporary fusion devices cannot achieve the plasma parameters of future machines, such as ITER [1] and DEMO [2], predictive modelling of thermal response and melt motion is crucial in guiding reactor designs and choices of safe operational space.
@@
The physics model and treatment of plasma-related effects have been guided by empirical evidence from a series of dedicated tokamak experiments [8, 9, 10, 11, 12, 13]. The numerical implementation of the model in the eponymous code [4, 5, 6] has been correspondingly carried out in an additive fashion. Moreover, as validation against experiments proved successful, the use of the tool for predictive modelling for fusion reactors became essential. Such applications pose additional challenges due to the necessity to simulate intricate PFC layouts as well as complex plasma-wetted geometries. Furthermore, simulations of at least an entire plasma-facing unit (monoblock, castellation) are necessary, where it is essential to resolve the thin localized melt pools on timescales that include long power loading combined with short quasi-periodic bursts due to edge localized modes (ELMs). The lack of a priori knowledge of the PFC geometries and wetting details as well as the spatiotemporal scales to be tackled inhibited the development of efficient numerical solutions. The original implementation of the MEMOS-U physics model was computationally inefficient and lacked flexibility in simulations of complex domains and complicated scenarios. This served as the primary incentive behind the development of a new implementation of the physics model in a modern computational tool - MEMENTO (MEtallic Melt Evolution in Next-step TOkamaks).
In this work, an in-depth overview of MEMENTO is presented, including a detailed description of the adopted numerical schemes. The MEMENTO code relies on AMReX, a software framework that provides functionality for massively parallel codes and adaptive meshing techniques [22, 23]. Numerical benchmarks of the solvers of MEMENTO have been presented in a previous publication [24]. Here, the code structure is discussed and a simulation flowchart is provided together with distinctive examples of the MEMENTO capabilities concerning the modelling of fusion-relevant melting events.
@@ 2 Model equations
In what follows, for the sake of completeness, the set of equations is recapped, while all the physical quantities and thermophysical properties involved are introduced:
@@
In the above, the liquid column height is defined as , with and denoting the evolving surface of the liquid-solid interface and the evolving metal-plasma interface (free surface) respectively, as sketched in Fig.1. The SW equations (1)-(2) and their BCs (1a), (2a) describe the evolution of the film thickness and the depth-averaged velocity ; is the rate of change of due to vaporization, is a unit vector tangential to the free surface, the bulk current density whose depth-averaged value is defined as
the external magnetic field, the mass density, the surface tension, the drag force and the ambient pressure applied by the plasma and vapour at the surface. In the heat equation (3) and its BC (3a); is the specific isobaric heat capacity, the thermal conductivity, the volumetric heat sources, any bound of the computational domain for which Eq.(3) is solved, the vector unit normal and the surface heat flux. In the EM equations (4)-(5) and their BCs (4a)-(4c) which enforce the conservation of charge; is the electrical resistivity, an auxiliary potential and the emitted current density. Three surface boundaries are distinguished; is the free surface with the unit vector normal , is any electrically insulated surface with the unit vector normal and is any grounded surface.
@@ 3 Code structure
MEMENTO relies on the AMReX framework [22, 23] to generate and maintain an adaptive non-uniform grid. AMReX is a C++ software with a Fortran interface that provides functionality for solving partial differential equations with block-structured and adaptive mesh refinement algorithms. The non-uniform grid allows MEMENTO to perform highly resolved calculations near the free surface and to save computational time by introducing a coarser grid far from it. Each distinct grid resolution defines an AMReX level, with the free surface always corresponding to the finest level. The coarsest level is always designated as “level 0”. The terms higher and finer will be used interchangeably when referring to AMReX levels. The ratio between grid resolutions of two different levels is referred to as the refinement ratio; it is defined in the input and has to be a power of two in all MEMENTO simulations. Fig.2 shows an example of the cross-section of a simulated sample and the employed three-level grid. Owing to the adaptive nature of the mesh, the grid has changed between the 5 and 11 ms instants due to the build-up of a strong surface deformation. To complement the non-uniform spatial mesh, sub-cycling in time can be enabled so that higher levels are advanced with a smaller time step. The ratio between the time steps of different levels is equal to their refinement ratio. In particular, with sub-cycling enabled, for every update of a coarser level, finer levels will undergo as many more updates as specified by the refinement ratio between them. For example, if there are two levels with a refinement ratio of 4, when advancing the coarse level in time once, the fine level is advanced four times with a 4 times smaller time step.
@@ 4.1 Implementation
The parameters and are controlled by the input and serve as ad-hoc solutions to possible numerical issues by capping the viscous damping and the Marangoni drive to a minimum effective fluid height. The designation refers to the partially updated height obtained by solving for the position of the free surface in time step but not for the solid-liquid interface, i.e. . It is noted that the explicit formulation of the surface tension introduces a stringent stability criterion for the time step that scales poorly with the tangential discretizations and [30]. Moreover, the Lorentz force, , is a variable that has to be passed from the electrostatics solver with the following interpolations employed
@@ 5 Heat transfer solver
The velocity field employed by the heat solver, , is a 3D vector, whereas the depth-averaged velocity from the SW solver is a 2D vector. Reconstruction of from is based on consistent with the SW assumption of . For ease of notation, in the following, will be simply expressed as .
@@ 5.2.1 Implementation
The time step indicator ∗ has the same meaning as the one discussed in the explicit heat solver, see section 5.1.1.
@@ 7 MEMENTO simulations
Possible simulated geometries are categorized in two groups; slab and complex. In slab geometries the initial free surface lies at a constant height, , while in complex geometries the free surface is a function of position, . Examples of complex geometries include, but are not limited to, free surfaces that are initially sloped, curved or step-like. The free surface is wetted by default, whereas the wetting of other surface boundaries is controlled by the input. Cooling is possible from all surfaces, but mass-flux due to vaporization is only defined for the free surface. The sample can include a heat sink in the solid which emulates a cooling pipe, as the one present in ITER monoblocks. Finally, it is possible to include regions where the thermal and electrical conductivities are set to zero, to emulate gaps in the domain.
@@ 8 Re-gridding, re-evaluation and synchronization
Re-gridding refers to the modification of the grid used by the heat and electrostatics solvers to comply with the deformation of the free surface. The goal is to keep the finely resolved region to the minimum size required at each time-instant in order to reduce the computational cost. In the current version of MEMENTO, there are three defined criteria that can trigger re-gridding to a higher or lower level. Those are the distance of a point from the free surface, the surface deformation rate and the material phase at a certain point. The values of the distance from the free surface and the deformation rate that trigger re-gridding are defined in the input. Note that the surface deformation rate is a function of and not . Hence, if the value triggers re-gridding then all the points for which are re-gridded. The third criterion sets all the points in the liquid phase at the highest level, thus the re-gridding is triggered as the melt front propagates. By default these criteria are checked at every time step while different frequencies can be chosen via input. The default option is to check for re-gridding at every time step. In Fig.2, at 5 ms re-gridding has been triggered only due to the distance from the free surface while at 11 ms also the surface deformation rate triggers re-gridding.
@@ 9 Solver benchmarking and comparison with experiments
Before illustrating the MEMENTO capabilities through the simulation of a tokamak experiment, the code benchmarking detailed in Ref.[24] will be summarized. The fluid shallow water solver has been compared with the augmented solver of the GeoClaw software [32, 33, 34, 35]. The accuracy of the explicit heat solver has been checked against analytical solutions of a two-region Neumann problem [36]. The implicit heat solver has been benchmarked utilizing the explicit heat solver as reference code. Furthermore, the MEMENTO implementation of a cooling pipe, based on simplifications suggested in Ref.[37], has been compared to the results of ANSYS simulations [38] and the results of simulations with a dedicated solver [39]. Finally, the electrostatics solver has been validated against simulations with the COMSOL Multiphysics® software [24]. The numerical inaccuracies were always found to be negligible compared to typical uncertainties of the MEMENTO input, primarily the experimental heat fluxes which often feature uncertainties in excess of 20 [3, 5, 6, 13].
@@ 10 Example simulation of a tokamak experiment
Given the above, the simulation of a particular recent tokamak experiment was selected in order to demonstrate the capabilities of MEMENTO. This experiment, featuring several momentum sources, a non-trivial initial free surface, a non-slab sample geometry and a particular plasma wetting pattern, exemplifies a complex scenario which MEMENTO has been designed to model.
@@ 10.1 Experimental scenario
The exposure of an iridium (Ir) sample to the ASDEX Upgrade (AUG) tokamak discharges will be modelled. The main physics aspects have been reported in Ref. [13]. In what follows, the focus lies on numerical considerations concerning the simulation set-up and post-processing.
@@ 10.3 Simulation outcome
Finally, the deformation profile is a key output of the simulation, since it can be directly compared to the experiment. Such a comparison is presented in Fig.10 revealing that MEMENTO is capable to reproduce the post-mortem evidence both qualitatively and quantitatively. Note that the shift between the MEMENTO deformation profile and the observations is well within the experimental uncertainties on the position of the heat-flux on the tile ( cm). For more detailed discussions on the melt dynamics and the final deformation, the reader is addressed to Ref.[13]. Once again, it is pointed out that deformation is accumulated from all discrete melting events and that resolidification takes place in between melting.
@@ 11 Summary and discussion
The status of MEMENTO’s solvers will be briefly recapped to point out possible future updates. Furthermore, inherent limitations of the code in specific fusion-relevant applications due to the underlying shallow water assumption, will be discussed.
@@
The fluid solver provides the solution to the shallow waters equations. Depending on the scenario, to address stability issues inherent to the shallow water model, it might require truncation of the melt depth column to some minimal values, set by the input. The introduction of such melt depth thresholds can be entirely avoided or set to negligible values by including the surface tension in the simulations; an optimal choice both from a physics and a numerical point of view. However, in scenarios where the melt pool is created at or moved to an edge, the shallow waters approximation cannot capture the 3D nature of the resulting pool curvature and the details of the corner wetting, which is a limitation of the physics model rather than of the implementation. In case the scenario features an insignificant forcing of the pool off the edge, then an impermeable wall can provide an adequate solution. Such a solution was employed in the modelling of a recent AUG melting experiment, where all features of the experimental surface deformation were successfully recovered in the MEMENTO runs [14]. Future optimization of this module primarily concerns the surface tension implementation. In particular, other numerical schemes will be investigated in an attempt to improve the flexibility with respect to the choice of grid resolution.
@@
Finally, it is stressed again that the shallow water approximation constitutes the major reason for MEMENTO’s ability to reliably model fusion-relevant melting scenarios where meter-long plasma-wetted areas must be simulated while resolving micrometer-deep melt pools. Simultaneously, MEMENTO’s implementation of shallow water averaging imposes constraints on certain types of scenarios. There are two associated limitations. The first limitation concerns cases where the free surface includes a curvature or an angle such that the local surface normal vectors (at positions where melting is realized) form an angle which is near or above degrees. This might occur because such features are either inherent to the initial free surface, as in Fig.11a, or generated when excessive melting and forcing result in drastic material excavations, as in Fig.11b. The second limitation concerns the fact that each pair of the tangential coordinates, i.e. the and directions for MEMENTO as illustrated in Fig.1, should uniquely specify one value of the free surface position . This means that wave breaking phenomena, as sketched in Fig.11c, cannot be modelled.
Since the shallow water originating limitations mainly concern the aforementioned two cases, it can be stated that the majority of complex fusion-relevant melting scenarios can be tackled with the MEMENTO simulations. Moreover, in order to facilitate the self-consistent modelling of small-scale melt dynamics and melt flow stability in fusion reactors, the MEMENTO simulation outcome can be utilized in specialized ANSYS CFD set-ups. Such a work-flow enables the faithful reproduction of the macroscopic picture (large wetted areas, three-dimensional heat transfer, long time thermal response, large scale melt dynamics) responsible for the eventual melt pool depths and velocities , while allowing for full detail simulations of a small domain of interest where and are imposed through boundary and initial conditions [42]. After the necessary tests against experimental evidence from melting events in the JET tokamak [42], this work-flow has been employed for predictive modelling of melt splashing in ITER [43] and will be utilized in predictive studies for DEMO.
@@ Appendix A Simulation flowchart
The flowchart of a MEMENTO simulation is shown in Fig.1. In short, a simulation starts by reading the input and performing the necessary initializations. The free surface is generated at the specified position, the initial temperature is set, the melt thickness is initialized to zero and the material properties are calculated. If sub-cycling is not enabled, the fluid solver is called first, followed by the heat solver and the electrostatics solver. If sub-cycling is enabled, the order of operations changes. The heat equation is first solved on all levels except the highest. Then, the shallow waters, heat and electrostatics solvers are called in sequence and repeatedly according to the time step division by the refinement ratio. Once all solvers are updated, the code checks an input-defined frequency to determine whether output is to be generated in the current time step. Finally, the termination conditions are checked to see whether the simulation has reached its end. If not, time is incremented and the calculations are repeated.